OpenStudy (anonymous):

A pile driver drives a post 27 inches into the ground on its first hit. Each additional hit drives the post 2/3 the distance of the prior hit. Find the total distance the post has been driven after 5 hits.

OpenStudy (imstuck):

The first hit goes in 27 inches. The second hit goes in\[27+\frac{ 2 }{3 }(27)=36\]The third hit goes in \[36+\frac{ 2 }{ 3 }(36)=54\]

OpenStudy (imstuck):

the fourth hit goes in \[54+\frac{ 2 }{3 }(54)=72\]and the last hit goes in\[72+\frac{ 2 }{ 3 }(72)=96\]

OpenStudy (imstuck):

in the end it is 96 inches down.

OpenStudy (anonymous):

Thank you!